OpenStudy (anonymous):

Having the polar and parametric equations of a cardioid : rho=2*cos(phi) + 3 x=rho*cos(phi) y=rho*sin(phi) find a way of rotating the cardioid 180° and 270° anticlockwise , modifing the original equations. Help me please!

OpenStudy (anonymous):

Yes, its hard :(

OpenStudy (anonymous):

@Zarkon

OpenStudy (sasogeek):

24 days seems like a long wait lol :) you have patience :)

OpenStudy (anonymous):

switching x and y for 270 would work by sheer luck?

OpenStudy (kinggeorge):

I believe that using \(\varphi+\theta\) where \(\theta\) is 180\(^\circ\) or 270\(^\circ\) counterclockwise would work.

OpenStudy (kinggeorge):

And \(\theta\) must be in radians.

OpenStudy (anonymous):

I know if I change rho=2*cos (phi) +3 to rho=2*sin(phi) + 3 , it turns 90° anticlockwise, I will check if rho=2*sin(phi + 3/2pi) + 3 works...

OpenStudy (kinggeorge):

After checking it myself, I think you need to use pi/2 instead of 3pi/2.

OpenStudy (anonymous):

I will check ;)

OpenStudy (anonymous):

THAKS TO EVERYONE!!

OpenStudy (anonymous):

THANKS JAJA